Let us suppose your company has the option to produce a component in-house for $20 per units, but they would have to invest in the machineries which costs $100000. Alternatively, they could buy it from a third party, which is charging $38 per unit.

What is the minimum quantity of components to make the investment in the machinery worthwhile?

**The answer is NOT 5555**
Answer #1

Answer:

  • Lets suppose X be the no. of units beyond which investing in machinery will be worthwhile. Then at this X level of production, production cost from third party and in-house will be equal.
  • Cost from third party = 38*X
  • Cost from in house = 100000 + (20*X)
  • So 38X = 100000+20X
  • On solving : X = 5555.55
  • So if the production is 5556 then it will be worthwhile investing in In-House.